As a member of the Finance Department of Ranch​ Manufacturing, your supervisor has asked you to compute the appropriate discount rate to use when evaluating the purchase of new packaging equipment for the plant. Under the assumption that the​ firm's present capital structure reflects the appropriate mix of capital sources for the​ firm, you have determined the market value of the​ firm's capital structure as​ follows: To finance the​ purchase, Ranch Manufacturing will sell 10​-year bonds paying interest at a rate of 7.2 percent per year​ (with semiannual​ payment) at the market price of ​$1 comma 069. Preferred stock paying a ​$1.99 dividend can be sold for ​$24.64. Common stock for Ranch Manufacturing is currently selling for ​$54.11 per share and the firm paid a ​$3.01 dividend last year. Dividends are expected to continue growing at a rate of 5.1 percent per year into the indefinite future. If the​ firm's tax rate is 30 ​percent, what discount rate should you use to evaluate the equipment​ purchase?

Bonds- $4,200,000 market value

preferred stock- $2,100,000 market value

common stock- $5,500,000 market value
Answer #1

Discount rate = WACC = 8.1%

1: After tax cost of debt

Using financial calculator
Input: FV=1000, PV=-1069, N=10*2=20, PMT=7.2%*1000/2 = 36

Solve for I/Y as 3.13059

YTM= I/Y*2 = 6.26%

After tax cost of debt = 6.26%*(1-30%) = 4.38%

2: Cost of Preferred stock= Annual dividend/price = 1.99/24.64 = 8.08%

3: cost of common equity = D1/Price+g

= 3.01*105.1%/54.11+5.1%

= 10.95%

4: WACC = sum of cost*weight

= 4.38%*42/(42+21+55)+8.08%*21/(42+21+55)+10.95%*55/(42+21+55)

=8.1%